Add the following vectors graphically. Use scale 1cm = 1m.


13. Two ropes in a vertical plane exert equal magnitude forces on a hanging weight but pull with

an angle of 86o between them. What pull does each one exerts if their resultant pull is 372 N

directly upward? Use component method. (N or Newton is unit of force, a vector quantity)


1
Expert's answer
2021-10-05T10:06:53-0400


Find the magnitude of the vertical component of either force:


"T_y=T\cos\frac{86°}2."

The resultant tension of 372 N, by the component method, is


"372=T_y+T_y=2T_y,\\\space\\
372=2T\cos\frac{86°}2,\\\space\\
T=\frac{372}{2\cos43°}=254\text{ N}."